OpenStudy (anonymous):

A probability experiment has four possible outcomes: e _1, e_2, e_3, e_4. The out come e_1 is four times as likely as each of the three remaining outcomes. Find the probability of e_1. what I got 4/7.

OpenStudy (nurali):

4p + p + p + p = 1 7p = 1 p = 1/7. P(e1) = 4p = 4(1/7) = 4/7

OpenStudy (anonymous):

Yes i got it correct thanks for the explanation.